OpenStudy (anonymous):

What is the value of the coefficient of "a" when the quadratic equation y = (5x − 2)(2x + 3) is written in standard form? −11 −6 7 10

OpenStudy (anonymous):

There is no "a" in the equation. Do you mean the formula: Ax + By = C

OpenStudy (anonymous):

yeah i guess soo.. cause thats the question.. idrk but your formula seems right

OpenStudy (anonymous):

Okay, I think you mean the quadratic formula. y = ax^2 + bx+ c y = (5x - 2)(2x + 3) y = 10x^2 + 15x - 4x - 6 y = 10x^2 + 11x - 6 Looking at that, what do you think is a?

OpenStudy (anonymous):

so a = 10?

OpenStudy (anonymous):

That's right! :)

OpenStudy (anonymous):

thanks :) you should help me on my other q's! id appreciate it

OpenStudy (anonymous):

Okay, just tag me in them. :)